A farmer has 40 metres of fencing with which he intends to make a rectangular enclosure for a few sheep and their lambs. The farmer calculates that he requires 96m²of grass in the sheep pen. By forming a quadratic equation find the dimensions of the pen
Answer by stanbon(75887)

A farmer has 40 metres of fencing with which he intends to make a rectangular enclosure for a few sheep and their lambs. The farmer calculates that he requires 96m²of grass in the sheep pen. By forming a quadratic equation find the dimensions of the pen
-----
Perimeter = 2(L + W) = 40 metres
----
Area = L*W = 96 m^2
----
Solve the 1st equation for "L":
L + W = 20
L = 20-W
----
Sustitute into the 2nd equation and solve for "W":
(20-W)W = 96
20W-W^2 = 96
W^2 - 20W + 96 = 0
----
(W -16)(W-4) = 0
Width = 16 m or Width = 4 m
---
If W = 16, then L = 4
If W = 4, then L = 16
=========================
Cheers,
Stan H.
=========================